OpenStudy (anonymous):

What addition fact can help you find 14-4?

OpenStudy (tester97):

7+9-2?

OpenStudy (anonymous):

You can use the fact that subtraction is the same as adding a negative. i.e.: \[a-b = a +(-b)\] In this case: \[14 - 4 = 14 +(-4) =10\]

OpenStudy (tester97):

or maybe 5+5 or 2+8 or 4+6